OpenStudy (anonymous):

lim h--> 0 (sin (pi/4+h) - sin pi/4)/ h

OpenStudy (phi):

do you know how to take the derivative of sin(x) ? or do you have to evaluate this limit without using the derivative?

OpenStudy (phi):

you could use the trig identity sin(A+B)= sin(A) cos(B) + cos(A) sin(B) you also need to know lim h->0 cos(h)= 1 and lim h->0 sin(h) = h

OpenStudy (phi):

because your limit is the definition of the derivative of sin(x) evaluated at x=pi/4 you could just do d/dx sin(x) = cos(x) and then evaluate cos(pi/4)

OpenStudy (anonymous):

Use sin(A+B)=sinAcosB+cosAsinB (sin(pi/4+h) - sin(pi/4))/h =(sin(pi/4)cos(h)+cos(pi/4)sin(h)-sin(pi/4))/h =(sqrt2/2)cos(h) + (sqrt(2)/2)*sin(h)-(sqrt2/2)/h this is where im stuck

OpenStudy (anonymous):

the answer should be 1/sqrt2

OpenStudy (phi):

you also need to know lim h->0 cos(h)= 1 and lim h->0 sin(h) = h

OpenStudy (anonymous):

but idk what to do next after (sqrt2/2)cos(h) + (sqrt(2)/2)*sin(h)-(sqrt2/2)/h

OpenStudy (phi):

cos(h) -->1 and sin(h)-->h so you have \[ \frac{(\sqrt{2}/2)1 + (\sqrt{2}/2)h-(\sqrt{2}/2)}{h} \]

OpenStudy (anonymous):

yes

OpenStudy (phi):

do you see sqr(2)/2 - sqr(2)/2 =0 leaving \[ \frac{(\sqrt{2}/2)h}{ h} \]

OpenStudy (anonymous):

yes i understand that then you cross/divide off the H which will give 1

OpenStudy (phi):

leaving \[ \frac{\sqrt{2}}{2} = \frac{1}{\sqrt{2}} \]

OpenStudy (anonymous):

omg thank you som much =)

OpenStudy (phi):

yw